OpenStudy (baldymcgee6):

The volume of revolution generated by rotating the function y = sin(2x)sqrt(cos(2x)) about the x-axis on the intevral [0, b] is pi/6 units^3, where 0

OpenStudy (amistre64):

ok, what do we have here

OpenStudy (baldymcgee6):

they gave us the volume, we need to find the upper limit somehow!

OpenStudy (amistre64):

the concept is simple enough .... F(b) - F(a) = given volume soo F(b) = given + F(a) and invert F(b)

OpenStudy (baldymcgee6):

right...? i'll take your word for it

OpenStudy (amistre64):

youre a trusting soul lol

OpenStudy (amistre64):

our function IS our radius, so go ahead and square it out

OpenStudy (baldymcgee6):

(sin(2x))^2*cos(2x)

OpenStudy (amistre64):

good, and if youve done enough of these, that should look rather familiar

OpenStudy (amistre64):

cos is the deriative of sin; so all we are missing is a 2 from that 2x

OpenStudy (amistre64):

and maybe a 3 from the derivative of ^3

OpenStudy (amistre64):

clean it up if you want with a usub; u = sin(2x) du = 2 cos(2x) dx du/2 = cos(2x) dx

OpenStudy (amistre64):

\[pi\int\frac{u^2 \ du}{2}\]

OpenStudy (baldymcgee6):

ahh i see. but then how do we find the upper limit?

OpenStudy (amistre64):

well\[\int_{a}^{b} f(x)\ dx = F(b)-F(a)=given\ value\]

OpenStudy (amistre64):

solve for b

OpenStudy (amistre64):

pi [sin(2x)]^3 ----------- ; at 0 this is 0 so that simply leaves us with 6 pi sin^3(2b)/6 = pi/6

OpenStudy (amistre64):

when we compare it all up we are left with sin^3(2b) = 1

OpenStudy (amistre64):

since sin(pi/2) = 1 2b = pi/2 b = pi/4

OpenStudy (baldymcgee6):

you are amazing

OpenStudy (amistre64):

nah, just lucky; i forgot the ^3 but that just hits the 1 and becomes pointless

OpenStudy (baldymcgee6):

so the answer is 0.79?

OpenStudy (amistre64):

the answer is pi/4 if you have to approx it; then by all means ... approx away

OpenStudy (baldymcgee6):

yeah, says to the nearest hundredth.. thanks so much

OpenStudy (amistre64):

your welcome

OpenStudy (amistre64):

but you see how the concept panned out right?

OpenStudy (baldymcgee6):

yeah, i never thought of it like that

OpenStudy (amistre64):

we left b alone till the end of it and it worked out for us :)

OpenStudy (baldymcgee6):

are you a teacher?

OpenStudy (amistre64):

not that im aware of :)

OpenStudy (baldymcgee6):

haha, well nonetheless, you rock!